Please help me solve this word problem:
The smaller of two #'s is 7 less than the Larger. If the smaller is subtracted from 6 times the Larger, the result is 72. Find the #'s.
I tried:
x= Larger #
x-7= Smaller #
6x-x-7=72
5x-7=72
5x=79 ??

Answer by stanbon(75887)

Please help me solve this word problem:
The smaller of two #'s is 7 less than the Larger. If the smaller is subtracted from 6 times the Larger, the result is 72. Find the #'s.
I tried:
x= Larger #
x-7= Smaller #
6x-(x-7)=72
5x+7=72
5x=65
x = 13
------------
x-7 = 6
-------------
Cheers,
Stan H.
-----------